# Multi-modal Evidential Fusion Network for Trusted PET/CT Tumor Segmentation

### v1

@@ Abstract
Accurate segmentation of tumors in PET/CT images is important in computer-aided diagnosis and treatment of cancer. The key issue of such a segmentation problem lies in the effective integration of complementary information from PET and CT images. However, the quality of PET and CT images varies widely in clinical settings, which leads to uncertainty in the modality information extracted by networks. To take the uncertainty into account in multi-modal information fusion, this paper proposes a novel Multi-modal Evidential Fusion Network (MEFN) comprising a Cross-Modal Feature Learning (CFL) module and a Multi-modal Trusted Fusion (MTF) module. The CFL module reduces the domain gap upon modality conversion and highlights common tumor features, thereby alleviating the needs of the segmentation module to handle modality specificity. The MTF module utilizes mutual attention mechanisms and an uncertainty calibrator to fuse modality features based on modality uncertainty and then fuse the segmentation results under the guidance of Dempster-Shafer Theory. Besides, a new uncertainty perceptual loss is introduced to force the model focusing on uncertain features and hence improve its ability to extract trusted modality information. Extensive comparative experiments are conducted on two publicly available PET/CT datasets to evaluate the performance of our proposed method whose results demonstrate that our MEFN significantly outperforms state-of-the-art methods with improvements of 2.15 and 3.23 in DSC scores on the AutoPET dataset and the Hecktor dataset, respectively. More importantly, our model can provide radiologists with credible uncertainty of the segmentation results for their decision in accepting or rejecting the automatic segmentation results, which is particularly important for clinical applications. Our code will be available at https://github.com/QPaws/MEFN.
@@
Tumor segmentation , PET/CT , Multimodal fusion Network , Dempster-Shafer theory , GAN1 Introduction
@@
PET (Positron Emission Tomography) and CT (Computed Tomography) serve as two major medical imaging techniques that play crucial roles in cancer diagnosis and treatment. CT images provide high-resolution anatomical information for tumor localization and edge detection and hence many researchers in recent years (Li et al., 2018; Yu et al., 2019) developed algorithms for the segmentation of tumor regions in the liver from CT images. However, they may struggle to precisely identify tumor boundaries due to density differences. PET images, on the other hand, measure the level of metabolic activity which reflects the functional status of tissues and cells and offers advantages for disease diagnosis and monitoring. Nevertheless, their lower resolution can lead to relatively blurred tumor boundaries and anatomical structures, which makes it difficult to meet the demands for segmentation accuracy. The complementarity of the CT and PET pushes the simultaneous examination with these two imaging modalities, which urges the development of joint tumor segmentation algorithms from PET/CT multi-modal images (Islam et al., 2023).
@@
In radiology, tumor segmentation plays a critical role in accurate diagnosis and treatment. The most critical challenge in multimodal segmentation is how to fuse the information from multiple modalities. Existing multimodal fusion methods can be categorized into three types: input-level fusion, feature-level fusion, and decision-level fusion. In input-level fusion, PET and CT images are typically fused by concatenation (Zhang et al., 2015; Chen et al., 2018). However, PET/CT joint segmentation requires capturing complementary information between the two modalities but this fusion approach often struggles to explore nonlinear relationships between modalities, mak-eing it difficult to fully utilize the inter-modality information. In feature-level fusion, high-dimensional features from different modalities are typically extracted separately and then combined together. However, simple feature fusion (Zhao et al., 2018) will introduce redundant or conflicting features, which can degrade network performance. Additionally, unlike MR/CT, there are significant feature differences between PET and CT modalities, which further complicates feature fusion process. In deci-sion-level fusion, researchers often use different segmentation networks to segment tumors in individual modalities separately (Nie et al., 2016). The segmentation results of each modality are then fused using specific fusion strategies (Rokach, 2010) to obtain the final segmentation result. However, in practical clinical applications, the quality of multi-modal data is often unstable (e.g., some modalities may be corrupted), so the quality of multi-modal inputs should be quantitatively measured in some way to provide doctors with predicting confidence, which is particularly important when deploying multi-modal segmentation models for clinical tasks (Ma et al., 2023).
With the development of deep learning, more and more deep neural networks (especially the convolutional neural networks (CNNs)) have been developed for segmenting tumors from medical images of different modalities. For deep segmentation models, the initialization of weights in neural networks significantly impacts their optimization. For instance, random initialization often leads to gradient vanishing, gradient explosion, or entrapment in saddle points and local optima during the optimization process, which renders the network difficult to train. To address these issues, authors in (Glorot and Bengio, 2010) took into account the variances of inputs and outputs and proposed the Xavier initialization method. This method maintains the gradient magnitudes of all layers approximately equal, which can help mitigate the problems of gradient vanishing or explosion and makes it suitable for training deep neural networks. However, for certain specific neural network architectures or activation functions (such as ReLU), gradient vanishing still persists. The authors in (He et al., 2015) devised the Kaiming initialization specifically for the ReLU activation function, which not only effectively avoids the issue of gradient vanishing and enhances the training efficiency but also improves performance of the network. Nonetheless, it is still not applicable to all activation functions. The orthogonal initialization method is another weights initialization strategy proposed in (Saxe et al., 2013) where weights were initialized by generating orthogonal matrices. This strategy can prevent mutual influence between weight matrices, thus aiding in reducing overfitting and improving the network’s generalization ability. However, this algorithm has a higher complexity and is not suitable for all neural network structures. Additionally, the aforementioned initialization methods are based on generic neural networks rather than a specific task. Therefore, in scenarios with high task difficulty (such as PET/CT segmentation), these methods still exhibit suboptimal performance.
@@
This paper aims to address the aforementioned issues and proposes a Multi-modal Evidential Fusion Network (MEFN) to achieve more accurate and reliable PET/CT multi-modal tumor segmentation. Specifically, this network comprises two main modules: the Cross-Modal Feature Learning (CFL) module for modality transfer and the Multi-modal Trusted Fusion (MTF) module for tumor segmentation. In CFL, we use GAN to enable back-and-forth transformation between the two modalities and set additional decoupling branches for the network to extract tumor features. The intuition behind this is that during the modality transformation process, the network can simultaneously learn the content features of both modalities. The proposed Multi-scale Tumor Guided Attention (MTGA) can assist the network in focusing more on extracting features from the tumor area and hence provides useful tumor information for subsequent segmentation. Then, the trained network parameters in CFL are transferred to the segmentation backbone of MTF and shared as the initial weights of the network during training. Such a task-driven weights initialization strategy can effectively improve the convergence speed, alleviate the aforementioned training challenges and achieve better performance. In MTF, in addition to the segmentation backbone, we also introduce the Dual-attention Feature Calibrating (DFC) module and adopt the Dempster-Shafer Theory Based Trusted Fusion (DBTF) method. These two methods respectively address feature-level fusion and decision-level fusion. Unlike previous fusion methods, our feature-level fusion takes into account the importance of weight allocated for different modalities and calibrates the semantic features extracted from each modality based on their confidence to minimize the impact of redundant and conflicting features on the network. The DBTF method, based on the Dempster-Shafer theory, performs decision-level fusion by considering the quality differences between different modalities and making fusion decisions based on uncertainty, thus fusing the segmentation results of different modalities both accurately and reliably.
In summary, this work makes the following three main contributions:
@@
A novel task-driven initialization method named CFL is proposed to alleviate the impact of improper weights initialization on network optimization.
@@ 2.1 Generative adversarial model
Generative Adversarial Networks (GANs) are based on zero-sum games in game theory and consist of a generator and a discriminator (Goodfellow et al., 2020). The generator creates data by learning its original distribution while the discriminator differentiates between real and generated data and hence guides the generator toward more realistic outputs. GAN is often used for data reconstruction (Wolterink et al., 2017; Wang et al., 2018), synthetic data generation (Bowles et al., 2018) and converting data between modalities (Ben-Cohen et al., 2019). For example, upon incorporating tumor segmentation labels into the generator, authors in (Bi et al., 2017) attempted to synthesize PET from CT using a conditional GAN (cGAN). However, this method is impractical for us as segmentation labels are the target predictions in our task. In (Zhang et al., 2022), the authors proposed Modality-Specific Segmentation Networks with integrated cGAN to extract common features from PET and CT images for lung tumor segmentation. However, this method merges image synthesis with segmentation tasks, which makes the computational overhead high. In (Zhang et al., 2021), a Cross-modal Feature Transition module was introduced with cycleGAN to facilitate cross-modal knowledge transfer, which aims to improve segmentation by using multi-modal information. This method, however, doesn’t focus enough on the modal differences crucial for segmentation in tumor areas. Motivated by the aforementioned method, we propose the Cross-Modal Feature Learning (CFL) module where the Multi-scale Tumor Guided Attention (MTGA) mechanism is introduced to assist concentrating on tumor areas while learning cross-modal features and reducing modal differences, thus enabling the network to learn the task-driven initial weights that are helpful for multi-modal tumor segmentation.
@@ 2.2 Uncertainty estimation
In clinical practice, AI often lacks sufficient decision-making information, which makes it crucial to quantify uncertainty for safe application (Begoli et al., 2019). Current research on uncertainty estimation includes Bayesian (Gal and Ghahramani, 2016), ensemble (Lakshminarayanan et al., 2017), deterministic (Van Amersfoort et al., 2020) and evidence-based methods (Huang et al., 2022). Bayesian methods in neural networks replace fixed parameters with probability distributions, leading to probabilistic output labels. For example, Monte Carlo dropout (MC dropout) is introduced as a Bayesian approximation in (Gal and Ghahramani, 2016), where dropout layers are used in multiple inference runs to mimic a Gaussian process. Although this strategy is computationally less demanding, it leads to inconsistent outputs (Kohl et al., 2018). Ensemble methods improve neural network classification by using model averaging for uncertainty estimation, involving training multiple models with random initialization and adversarial training (Kamnitsas et al., 2018; Lakshminarayanan et al., 2017). While this method doesn’t require altering the network architecture, it necessitates retraining models from scratch and hence incurs high computational costs for complex models. Deterministic-based methods, like Deep deterministic uncertainty (Mukhoti et al., 2021), have been extended to semantic segmentation using feature space density. However, they alter network structure and increase computational costs. A recent study (Huang et al., 2022) suggested using deep feature extraction and evidence layers for segmenting lymphoma in PET and CT scans. Rather than to obtain more robust segmentation by calibrating the uncertainty, these studies tried to to improve the performance of segmentation upon guiding the the uncertainty. They only utilized the generated uncertainty to evaluate the segmentation results but not to further optimize the network training. Instead, our approach introduces not only an Uncertainty Perceptual Loss for robust segmentation by forcing the network focusing on the prediction results of regions with large uncertainty during training but also an Uncertainty Calibrator (UC) to further refine network training by calibrating the extracted semantic information according to uncertainty.
@@ 2.3 Multi-modal fusion
Multi-modal medical image segmentation, essential for accurately identifying tumors, employs various fusion strategies: input-level, feature-level and decision-level fusion. A typical example of input-level fusion is the 3D-Inception ResNet proposed in (Qayyum et al., 2021) which combined PET and CT images in channel-wise manner. Some recent studies (Wang et al., 2021; Liu et al., 2021) employed attention mechanisms in their models to allocate varying weights to different modality spaces and achieved fairly better results. Although the input-level fusion preserves the original details of the image, it fails in utilizing of the interrelations between different modalities. Feature-level fusion trains networks with individual modal images and then fuses features of different modalities across layers for final segmentation. It can effectively integrate and utilize multi-modal images (Zhou et al., 2019). Most existing methods in literature fuse multi-modal features by directly combining features of each modality (Zhong et al., 2018), which potentially miss key information and introduce redundancies and hence hinder network performance. Ahmad et al. (2023) introduced the Anatomy Aware Tumor Segmentation Network (AATSN), which uses an encoder to extract CT anatomical and an encoder to extract PET metabolic features for a lightweight fusion attention decoder. In this way, the decoder can use CT for spatial and PET for metabolic reference and hence improves tumor segmentation. However, its lightweight attention mechanism might under-perform with complex data due to limited training parameters. Decision-level fusion employs separate networks for each modality image. Each network uses its modality specific information whose outputs are integrated for the final result. Conventional fusion strategies include averaging and majority voting (Rokach, 2010). Kamnitsas et al. (2018) used an averaging strategy where the authors trained three networks and averaged their confidence scores for voxel segmentation. The majority voting strategy assigns voxels the most frequent label from all networks. However, these methods often overlook the uncertainty in each modality and lead to potential errors in probabilistic fusion, which is fatal in clinical practice. Huang et al. (2021a) combined PET and CT results using a Dempster-Shafer Theory-based fusion layer in 3D UNet. But this method, limited to decision layer fusion, shows less efficiency in handling multi-modal information. To resolve this problem, this paper merges feature-level and decision-level fusion strategies with the Dual-attention Feature Calibrating (DF
C) module and Dempster-Shafer Theory Based Trusted Fusion (DBTF) method, respectively. DFC reduces feature redundancies and calibrates low-confidence outputs. It’s worth noting that, differing from Huang et al. (2021a), our DBTF fuses PET and CT data based on uncertainty and mutual information, which greatly improves multi-modal fusion efficiency.
@@ 3 Methodology
In this section, we commence by succinctly outlining the MEFN framework, before delving into a comprehensive exposition of its three pivotal components: the Cross-Modal Feature Learning (CFL) module, the Dual-attention Feature Calibrating (DFC) module and the Dempster-Shafer Theory Based Trusted Fusion (DBTF). The overall framework of MEFN is provided in Fig. 1.
@@ 3.1 Overview of MEFN
Defining CT image as and PET image as , the two inputs ( and ) are first fed into the CFL. Within the CFL, it orchestrates the generation of pseudo PET images and pseudo CT images for each modality through adversarial training. Concurrently, an auxiliary tumor decoder branch produces corresponding artificial tumor mask images . In the training phase, the network acquires modality-specific features of CT and PET along with general tumor features, thereby minimizing the disparities between modalities. Subsequently, the network parameters trained in the CFL are transferred to the segmentation backbone as the initial weights of the network, therefore accelerating the convergence speed and reducing the risk of falling into local minima under the task-driven initialization of weights. Next, the encoder and decoder features from the two segmentation backbones are fed into the DFC. Within the DFC, encoder and decoder features undergo two distinct fusion strategies: in terms of Cross Modality Attention (CMA) and in terms of Uncertainty Calibrator (UC), respectively. The former adaptively adjusts features extracted by the encoder through mutual attention, which significantly reduces feature redundancies and enhances fusion efficiency. The latter calibrates semantic features extracted by the network using uncertainty factors, which forces the network focusing on areas of high uncertainty within modalities and thus diminishes feature confliction between modalities and enhancing fusion accuracy. Features refined by CMA and UC are then input into the fusion decoder to obtain the fused segmentation results and after upsampling. Finally, segmentation results from the CT branch ( and ), PET branch ( and ), along with the merged segmentation results ( and ), are collectively processed in terms of DBTF for decision-level fusion. In DBTF, the segmentation results of each branch will dynamically adjust the fusion weight according to the uncertainty, making the fusion result more robust, then the final segmentation outcomes can be acquired as and . Although the network structures in CFL and MTF are not necessarily the same, the more network layers these two networks share, the richer the modality-specific features can be transferred from CFL to MTF. To this end, we have designed two modules with quite similar network structures and the only difference is the number of kernels set to the decoder’s output head. As shown in Fig. 1, the decoder’s output head in CFL () uses one convolutional kernels and employs a activation function, while the number of convolutional kernels in the decoder’s output head in MTF () corresponds to the number of segmentation classes and a activation function is utilized in this case. As can be seen, such an MTF network architecture could share the most network layers with the CFL and thus can take full advantage of the features learned from the CFL. The algorithm for the overall process of MEFN can be described as in Algorithm 1.

### v2

# Multi-modal Evidential Fusion Network for Trustworthy PET/CT Tumor Segmentation

@@ Abstract
Accurate tumor segmentation in PET/CT images is crucial for computer-aided cancer diagnosis and treatment. The primary challenge lies in effectively integrating the complementary information from PET and CT images. In clinical settings, the quality of PET and CT images often varies significantly, leading to uncertainty in the modality information extracted by networks. To address this challenge, we propose a novel Multi-modal Evidential Fusion Network (MEFN), which consists of two core stages: Cross-Modal Feature Learning (CFL) and Multi-modal Trustworthy Fusion (MTF). The CFL stage aligns features across different modalities and learns more robust feature representations, thereby alleviating the negative effects of domain gap. The MTF stage utilizes mutual attention mechanisms and an uncertainty calibrator to fuse modality features based on modality uncertainty and then fuse the segmentation results under the guidance of Dempster-Shafer Theory. Besides, a new uncertainty perceptual loss is introduced to force the model focusing on uncertain features and hence improve its ability to extract trusted modality information. Extensive comparative experiments are conducted on two publicly available PET/CT datasets to evaluate the performance of our proposed method whose results demonstrate that our MEFN significantly outperforms state-of-the-art methods with improvements of 3.10 and 3.23 in DSC scores on the AutoPET dataset and the Hecktor dataset, respectively. More importantly, our model can provide radiologists with credible uncertainty of the segmentation results for their decision in accepting or rejecting the automatic segmentation results, which is particularly important for clinical applications. Our code will be available at https://github.com/QPaws/MEFN.
@@
Tumor segmentation , PET/CT , Multimodal fusion Network , Dempster-Shafer theory , Generative Adversarial Networks1 Introduction
@@
PET (Positron Emission Tomography) and CT (Computed Tomography) serve as two major medical imaging techniques that play crucial roles in cancer diagnosis and treatment. CT images provide high-resolution anatomical information for tumor edge detection and hence many researchers in recent years [1, 2] developed algorithms for the segmentation of tumor regions in the liver from CT images. However, these methods often face challenges in accurately locating and delineating tumors due to tissue density variations, noise, and limited contrast, especially for small or subtle lesions. PET images, on the other hand, measure the level of metabolic activity which reflects the functional status of tissues and cells and offers advantages for disease diagnosis and monitoring. Nevertheless, their lower resolution can lead to relatively blurred tumor boundaries and anatomical structures, which makes it difficult to meet the demands for segmentation accuracy. The complementarity of the CT and PET pushes the simultaneous examination with these two imaging modalities, which urges the development of joint tumor segmentation algorithms from PET/CT multi-modal images [3].
@@
In radiology, tumor segmentation plays a critical role in accurate diagnosis and treatment. The most critical challenge in multimodal segmentation is how to fuse the information from multiple modalities. Existing multimodal fusion methods can be categorized into three types: input-level fusion, feature-level fusion, and decision-level fusion. In input-level fusion, PET and CT images are typically fused by concatenation [4, 5]. However, PET/CT joint segmentation requires capturing complementary information between the two modalities but this fusion approach often struggles to explore nonlinear relationships between modalities, mak-eing it difficult to fully utilize the inter-modality information. In feature-level fusion, high-dimensional features from different modalities are typically extracted separately and then combined together. However, simple feature fusion [6] will introduce redundant or conflicting features, which can degrade network performance. Besides, the existence of a domain gap and the significant feature differences between PET and CT modalities, unlike the relatively consistent features in MR/CT, further complicate the process of effectively fusing information from these two modalities within a network. In decision-level fusion, researchers often use different segmentation networks to segment tumors in individual modalities separately [7]. The segmentation results of each modality are then fused using specific fusion strategies to obtain the final segmentation result. However, in practical clinical applications, the quality of multi-modal data is often unstable (e.g., some modalities may be corrupted), so the quality of multi-modal inputs should be quantitatively measured in some way to provide doctors with predicting confidence, which is particularly important when deploying multi-modal segmentation models for clinical tasks [8].
@@
This paper aims to address the aforementioned issues and proposes a Multi-modal Evidential Fusion Network (MEFN) to achieve more accurate and reliable PET/CT multi-modal tumor segmentation. Specifically, this network comprises two training stage: the Cross-Modal Feature Learning (CFL) stage based on modality translation and multi-task learning for more robust representation learning and the Multi-modal Trustworthy Fusion (MTF) stage for multimodal tumor segmentation. These two stages are alternately trained to improve feature representation and segmentation accuracy. In CFL stage, we use Generative Adversarial Networks (GAN) to enable transformation between the two modalities and set additional tumor decoders to guide the network to focus on tumor features. Then, the trained network parameters in CFL are shared with the segmentation backbone of MTF and as the initial weights of the network during training. Such modality translation aligns features across different modalities into a shared feature space, enhancing the model’s ability to capture relevant features and improving segmentation performance with a more robust multimodal representation during the MTF stage. In MTF stage, in addition to the segmentation backbone, we also introduce the Dual-attention Feature Calibrating (DFC) module and adopt the Dempster-Shafer Theory-based Trustworthy Fusion (DTF) method. These two methods respectively address feature-level fusion and decision-level fusion, considered the uncertainty between different modalities and reduced the impact of redundant and conflicting features on the model.
In summary, this work makes the following main contributions:
@@
A GAN-based modality translation and multi-task constrained CFL stage is proposed to obtain more robust representation of the model to alleviate the impact of domain gap between different modalities.
@@ 2.1 Generative adversarial model
Generative Adversarial Networks (GAN) are based on zero-sum games in game theory and consist of a generator and a discriminator [9]. The generator creates data by learning its original distribution while the discriminator differentiates between real and generated data and hence guides the generator toward more realistic outputs. GAN is often used for data reconstruction [10, 11], synthetic data generation [12] and converting data between modalities [13]. For example, upon incorporating tumor segmentation labels into the generator, Bi et al. [14] attempted to synthesize PET from CT using a conditional GAN (cGAN). However, this method is impractical for us as segmentation labels are the target predictions in our task. Zhang et al. [15] proposed Modality-Specific Segmentation Networks with integrated cGAN to extract common features from PET and CT images for lung tumor segmentation. However, this method merges image synthesis with segmentation tasks, which makes the computational overhead high. Ding et al. [16] proposed a cross-modality transfer learning to allow knowledge learned from one imaging modality to be transferred to improve segmentation performance on another modality. This method, however, only considers one-way conversion. Motivated by the aforementioned method, we propose the Cross-Modal Feature Learning (CFL) stage where the Tumor Guided Attention (TGA) mechanism is introduced to help the model focus on tumor regions while simultaneously learning more robust representations and alleviating the impact of the domain gap.
@@ 2.2 Uncertainty estimation
In clinical practice, AI often lacks sufficient decision-making information, which makes it crucial to quantify uncertainty for safe application [17]. Current research on uncertainty estimation includes Bayesian [18], ensemble [19], deterministic [20] and evidence-based methods [21]. Bayesian methods in neural networks replace fixed parameters with probability distributions, leading to probabilistic output labels. For example, Monte Carlo dropout (MC dropout) is introduced as a Bayesian approximation [18], where dropout layers are used in multiple inference runs to mimic a Gaussian process. Although this strategy is computationally less demanding, it leads to inconsistent outputs [22]. Ensemble methods improve neural network classification by using model averaging for uncertainty estimation, involving training multiple models with random initialization and adversarial training [23, 19]. While this method doesn’t require altering the network architecture, it necessitates retraining models from scratch and hence incurs high computational costs for complex models. Deterministic-based methods, like Deep deterministic uncertainty [24], have been extended to semantic segmentation using feature space density. However, they alter network structure and increase computational costs. A recent study [21] suggested using deep feature extraction and evidence layers for segmenting lymphoma in PET and CT scans. Rather than to obtain more robust segmentation by calibrating the uncertainty, these studies tried to to improve the performance of segmentation upon guiding the the uncertainty. They only utilized the generated uncertainty to evaluate the segmentation results but not to further optimize the network training. Instead, our approach introduces not only an Uncertainty Perceptual Loss for robust segmentation by forcing the network focusing on the prediction results of regions with large uncertainty during training but also an Uncertainty Calibrator (UC) to further refine network training by calibrating the extracted semantic information according to uncertainty.
@@ 2.3 Multi-modal fusion
Multi-modal medical image segmentation, essential for accurately identifying tumors, employs various fusion strategies: input-level, feature-level and decision-level fusion. A typical example of input-level fusion is the 3D-Inception ResNet [25], which combined PET and CT images in channel-wise manner. Some recent studies [26, 27] employed attention mechanisms in their models to allocate varying weights to different modality spaces and achieved fairly better results. Although the input-level fusion preserves the original details of the image, it fails in utilizing of the interrelations between different modalities. Feature-level fusion trains networks with individual modal images and then fuses features of different modalities across layers for final segmentation. It can effectively integrate and utilize multi-modal images [28]. Most existing methods in literature fuse multi-modal features by directly combining features of each modality [29], which potentially miss key information and introduce redundancies and hence hinder network performance. Ahmad et al. [30] introduced the Anatomy Aware Tumor Segmentation Network (AATSN), which uses an encoder to extract CT anatomical and an encoder to extract PET metabolic features for a lightweight fusion attention decoder. In this way, the decoder can use CT for spatial and PET for metabolic reference and hence improves tumor segmentation. However, its lightweight attention mechanism might under-perform with complex data due to limited training parameters. Decision-level fusion employs separate networks for each modality image. Each network uses its modality specific information whose outputs are integrated for the final result. Conventional fusion strategies include averaging and majority voting. Kamnitsas et al. [23] used an averaging strategy where the authors trained three networks and averaged their confidence scores for voxel segmentation. The majority voting strategy assigns voxels the most frequent label from all networks. However, these methods often overlook the uncertainty in each modality and lead to potential errors in probabilistic fusion, which is fatal in clinical practice. Huang et al. [31] combined PET and CT results using a Dempster-Shafer Theory-based fusion layer in 3D UNet. But this method, limited to decision layer fusion, shows less efficiency in handling multi-modal information. To resolve this problem, this paper merges feature-level and decision-level fusion strategies with the Dual-attention Feature Calibrating (DFC) module and Dempster-Shafer Theory-based Trustworthy Fusion (DTF) method, respectively. DFC reduces feature redundancies and calibrates low-confidence outputs. It’s worth noting that, differing from Huang et al. [31], our DTF fuses PET and CT data based on uncertainty and mutual information, which greatly improves multi-modal fusion efficiency.
@@ 3 Methodology
In this section, we commence by succinctly outlining the MEFN framework, before delving into a comprehensive exposition of its three pivotal components: the Cross-Modal Feature Learning (CFL) stage, the Dual-attention Feature Calibrating (DFC) module and the Dempster-Shafer Theory Based Trustworthy Fusion (DTF). The overall framework of MEFN is provided in Fig. 1.
@@ 3.1 Overview of MEFN
Defining CT image as and PET image as , the two inputs ( and ) are first fed into the CFL. Within the CFL, it orchestrates the generation of pseudo PET images and pseudo CT images for each modality through adversarial training. Concurrently, an auxiliary tumor decoder branch produces corresponding artificial tumor mask images . In the training phase, the network acquires modality-specific features of CT and PET along with common tumor features, thereby learning more robust feature representations. Subsequently, the network parameters trained in the CFL are shared to the segmentation backbone as the initial weights of the network. Next, the encoder and decoder features from the two segmentation backbones are fed into the DFC. Within the DFC, encoder and decoder features undergo two distinct fusion strategies: in terms of Cross Modality Attention (CMA) and in terms of Uncertainty Calibrator (UC), respectively. The former adaptively adjusts features extracted by the encoder through mutual attention, which significantly reduces feature redundancies and enhances fusion efficiency. The latter calibrates semantic features extracted by the network using uncertainty factors, which forces the network focusing on areas of high uncertainty within modalities and thus diminishes feature confliction between modalities and enhancing fusion accuracy. Features refined by CMA and UC are then input into the fusion decoder to obtain the fused segmentation results and after upsampling. Finally, segmentation results from the CT branch ( and ), PET branch ( and ), along with the fused segmentation results ( and ), are collectively processed in terms of DTF for decision-level fusion. In DTF, the segmentation results of each branch will dynamically adjust the fusion weight according to the uncertainty, making the fusion result more robust, then the final segmentation outcomes can be acquired as and . Although the network structures in CFL and MTF are not necessarily the same, the more network layers these two networks share, the richer the modality-specific features can be transferred from CFL to MTF. To this end, we have designed two stage with quite similar network structures and the only difference is the number of kernels set to the decoder’s output head. As shown in Fig. 1, the decoder’s output head in CFL () uses one convolutional kernels and employs a activation function, while the number of convolutional kernels in the decoder’s output head in MTF () corresponds to the number of segmentation classes and a activation function is utilized in this case. As can be seen, such an MTF network architecture could share the most network layers with the CFL and thus can take full advantage of the features learned from the CFL.